# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ 3.1.1 Aligned and Anti-aligned Magnetic Moments
[…]
@@
[…]
Figure 9: The polar caps, i.e., instantaneous zones of open flux, of the two stars in a Mollweide projection in the case with $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ},\,\lambda=-1$ and $\Delta\Phi=0^{\circ}$(
@@ 3.1.2 Inclined Magnetic Moments
Inclined magnetic moments introduce significant richness into the solutions, breaking both the latitudinal symmetry and the reflection symmetry of the emergent Poynting flux across the axial plane that passes through the center of the system, perpendicular to the line connecting the two stars. Figure 8 depicts configurations where the stars have non-zero inclination angles for both the quasi-aligned ($\lambda>0$, first and second columns) and quasi-anti-aligned ($\lambda<0$, third and fourth columns) solutions. In the first ( $\uparrow$ $\uparrow$ ) and third columns ( $\uparrow$ $\downarrow$ ), a star has an inclination angle of $45^{\circ}$, whereas in the second ( $\uparrow$ $\uparrow$ ) and fourth columns ( $\uparrow$ $\downarrow$ ), it has an inclination angle of $90^{\circ}$. Notably, the second and fourth columns correspond to configurations that are mirror images of each other under a rotation by $\pi$ around the axis, defined by the line connecting the two stars. Additionally, in the corotating frame, the two stars appear to rotate because they are irrotational in the observer frame.
Unlike the fully aligned or anti-aligned cases (§3.1.1), the presence of inclined magnetic moments introduces a more pronounced impact of orbital motion on the magnetosphere’s structure. The resulting magnetospheres evolve dynamically, exhibiting periodic variations and eruptions of Poynting flux linked to the orbital phases. As the stars orbit each other, magnetic fields originating from one star are carried along with its motion, anchored on its surface. As a result, the system reorganizes during a single rotation (see the middle and bottom rows in Figure  5). Field lines can dynamically evolve throughout the inspiral, some may connect one star to the other, while others that initially link the two stars may break and reconfigure. As a result, field lines anchored to a star’s surface can transition between being open, closed, or connecting with the companion star.
Thus, the polar caps, defined as footpoints of field lines extending to infinity (i.e., open flux), evolve over time, as shown in Figure 9. This evolution follows the orbital periodicity, with the variability becoming increasingly pronounced as the separation between the stars decreases during inspiral. The polar cap regions are found by integrating magnetic field lines using a fourth-order Runge-Kutta method, classifying as open those that cross a cylinder of radius $25r_{\star}$ and height of $50r_{\star}$. In Figure 9, the polar caps of both stars are shown at two snapshots separated by half a rotation, for the case with $\lambda=-1$, $\alpha_{1}=0^{\circ}$ and $\alpha_{2}=45^{\circ}$ ( $\uparrow$ $\downarrow$ ).
@@
Figure 10 shows the flux distributions on a spherical surface with radius $r=13r_{\star}$ corresponding to the magnetospheric structures and snapshots shown in Figure 8. Consistent with the discussion above, unlike the zero-inclination cases discussed previously, the flux distribution in these cases is no longer symmetric between the northern and southern hemispheres. While the flux pattern may appear symmetric over a full orbital period, the instantaneous flux evolves throughout the inspiral, leading to unequal fluxes in the two hemispheres at any given moment. This asymmetry in the instantaneous and time-averaged Poynting flux pattern indicates the presence of EM forces and torques acting on the system, influencing the motion of the NSs and potentially affecting the post-merger remnant. This aspect is further explored in Section 3.2.
@@
The variability and dependence on magnetic configuration of the EM flux is also evident in Figure 12, which shows the total flux passing through the spherical surface with $r=13r_{\star}$ as a function of time. For $\lambda>0$ (left panel), the total flux increases when the stars have non-zero inclination angles and rises further as the angles increase. In contrast, for $\lambda<0$ (right panel), the total flux decreases with increasing inclination angles. Notably, when both stars have inclination angles of $90^{\circ}$ and $\lambda<0$, the total flux can drop to nearly zero at specific times during the evolution. This is illustrated in the bottom row of Figure  5, where two snapshots are shown, separated by an orbital phase of $\pi/2$. When the magnetic poles of the two stars with the same polarity face each other (left panel), field lines bend away, twist, and predominantly connect each star’s opposite magnetic poles, allowing a modest number of field lines to remain open. On the other hand, after a time evolution corresponding to the rotation of $\pi/2$ (right panel), the magnetic moments are oriented side by side, pointing in opposite directions. This configuration leads to even fewer open field lines and a significantly lower escaping flux compared to the first snapshot.
This behavior reflects the fact that the aligned and anti-aligned configurations represent two extreme cases of the angle between the magnetic moments, directly influencing the degree of connectivity between the stars and the amount of open flux $\Phi_{\rm open}$.
For completeness, it is worth noting how the above results change in the opposite limit of tidally locked (synchronously rotating) stars, where $\omega_{1}=\omega_{2}=\Omega$. The differences relative to the irrotational cases depend on the magnetic configuration. In the synchronized and aligned case, the overall magnetospheric structure and resulting Poynting flux would be nearly identical to the aligned irrotational configuration. This is because the relative orientation between the magnetic moments and the orbital motion remains constant, resulting in minimal time variation in the magnetosphere. In the synchronized and anti-aligned configuration, some magnetic field lines connect the two stars, reducing the number of open field lines available to transport the EM energy. This leads to a lower Poynting flux compared to the corresponding aligned case, in contrast to what was found in the irrotational scenario. However, for cases where at least one of the stars has an inclined magnetic moment, the differences are more pronounced. The synchronization prevents the time-varying reconfiguration of field lines seen in the irrotational cases. As a result, the Poynting flux evolution arises solely from the shrinking orbital separation, which increases the polar cap area and flux magnitude.
@@ 3.1.3 Unequal Magnetization Configurations
Figure 12: The total Poynting flux crossing a spherical surface with radius $13r_{\star}$ for various inclination angle combinations of the two stars, shown for cases with $\lambda=1$ (Left) and $\lambda=-1$ (Right).The dotted line, in both plots, is the orbital average Poynting flux $\overline{\langle L\rangle}$ derived from the scaling law of Eq. 15 for $|\lambda|=1$
@@
Figure 13: The magnetospheric structure at $\tau=-4.27$ ms, similar to Figure 8, is shown with increasing absolute values of the magnetic moment ratio from top to bottom. The left column illustrates the impact of the magnetic moment ratio in the anti-aligned case ($\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$) (
@@
The distribution of orbit-averaged emission in cases with $|\lambda|>1$ becomes increasingly concentrated around the equator, as shown in the left column of Figure 14. As $|\lambda|$ increases, the emission progressively resembles that of an isolated pulsar with the same inclination angle as the dominant star, similar to Fig. 2.
@@ 3.1.4 Anisotropic Emission Patterns
[…]
@@
In particular, in general, $50\%$ of the escaping EM energy is in $<50\%$ of the total sky area, while in a some configurations, it is directed toward $<15\%$ of the sky (see Figure 16).
In specific configurations, the emission anisotropy is more uniform, similar to an isolated pulsar. In these cases, the time-averaged energy distribution is independent of the azimuthal angle; thus potential observers located at different azimuths would receive similar amounts of fluence with the observer’s polar angle, $\zeta$ being the only determining parameter.
[…]
@@ 3.1.5 Scaling Laws for Poynting Luminosity and the Binary Braking Index
Simulating the evolution of the magnetosphere for a variety of configurations allows us to investigate how the escaping EM flux depends on the orbital frequency, $\Omega$, and consequently on time via Eq. 5. The simplest way to quantify this dependency is to assume a power-law relation between the luminosity and the orbital frequency analogous to an isolated pulsar, such that $\langle L\rangle\propto\Omega^{p}$, where $p$ is the power-law index. The classical dipolar isolated pulsar magnetosphere in either vacuum or FF regimes follows $p\equiv 4$ for the far-field, and it has often been assumed by heuristic arguments arguments that this law roughly (e.g. $p\sim 3.3-4.7$, see discussion) holds in the double NS case. By fitting this function to the evolving total Poynting flux from each BNS simulation, we obtain the distribution of the $p$ values, shown in Figure 17.
@@
Figure 17: Histogram of the power-law index $p$ obtained from fitting of the relation $\langle L\rangle\propto\Omega^{p}$ shown in gray for all data points and in red only for the peak values. The distribution of $p$ spans a wide range from $\sim 1$ to $>6$, in contrast to the $p\equiv 4$ value for an isolated dipolar pulsar.
@@
Notably, the parameter $p$ now spans a broad range of values spanning from 1 to more than 6. While there are many cases around $p=4$, the binary system shows cases where $p$ deviates significantly, either higher or lower, due to the complex interactions between the two stars’ magnetospheres. This surprisingly large deviation from the conventional dipole braking index $n\equiv p-1=3$ highlights the richness of near-field effects permitted in the double NS case that modify far fields and escape radiation.
@@
Figure 18: Examples fit of the Poynting flux luminosity with a power law $\langle L\rangle\propto\Omega^{p}$. Top row: (Left) The case with $\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$ and $\lambda=-1$. (Right) The case with $\alpha_{1}=45^{\circ},\alpha_{2}=0^{\circ}$, and $\lambda=10$ demonstrates that although the power law cannot describe the periodic fluctuations, it captures well the trend of the luminosity. Bottom row: (Left) The case with $\alpha_{1}=90^{\circ},\alpha_{2}=90^{\circ},\lambda=-1$ and $\Delta\Phi=0^{\circ}$ highlights that although the flux can vary significantly, dropping to very low values and affecting the flux dependence on $\Omega$ the power law still describes quite well the averaged trend. (Right)The case with $\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$ and $\lambda=100$, which consists of an example of not that accurate description of the Poynting flux luminosity’s evolution with the power law. In all the cases, the dotted line indicates the moving average for comparison.
@@
Figure  18 showcases characteristic examples of the Poynting flux luminosity fits. While the power-law model captures the gross secular orbit-averaged trend of the luminosity, it does not account for the periodic fluctuations around the fitted $L(\Omega)$ values. This is particularly evident for high inclination angles (bottom-left panel in Figure 18), where, as described in the previous section, the total flux fluctuates dramatically, occasionally dropping to near-zero levels. This extreme variability weakens the overall dependence of the flux on $\Omega$, causing the value of $p$ to decrease, often approaching values close to 1, as seen in Figure 17. Importantly, a fitted $p\sim 1$ does not necessarily imply an induced split-monopolar or wind configuration (e.g., A. K. Harding et al., 1999; N. Bucciantini et al., 2006). Instead, it reflects the presence of intense, episodic bursts of Poynting flux that produce similar luminosities from one orbit to the next.
@@
In some cases, the evolution of the Poynting flux appears to transition between two distinct regimes (bottom-right panel in Figure 18), resulting in a fitted $p$ that represents an average rather than a single consistent trend. For instance, in the configuration shown in the bottom-right panel, where both magnetic moments are aligned (to each other and to $\boldsymbol{\Omega}$), and one star’s field dominates, no field lines connect the two stars. At first, only the region of the open field lines of the primary star are influenced mostly by the secondary, with some bending around the companion and closing back onto the primary. This phase corresponds to the initial rise behavior in Poynting flux. However, as the inspiral progresses, the companion moves into the closed-field region of the primary, leading to a transition where more field lines open. Since no lines can directly connect the two stars in this scenario, these newly opened field lines increase the Poynting flux faster, modifying its overall evolution.
@@
Despite these variations, we find that in the majority of cases, the average Poynting flux evolution follows a single well-defined power-law trend over the fitting interval in our simulations. Instances where the flux transitions between two regimes, as seen in the bottom-right panel of Figure 18, remain relatively rare.
@@
Figure 19: Scatter plot showing the power $p$ derived from the power law fitting $\langle L\rangle\propto\Omega^{p}$ for the different simulated cases. On $x$ axis is $\lambda$ in logarithmic scale while the color and the size of each point indicate the two inclination angles $\alpha_{1}$ and $\alpha_{2}$ according to the legends. The circles are cases with $\lambda>0$, while the triangles are with $\lambda<0$. The different $p$ values for the same $\alpha_{1}$, $\alpha_{2}$ and $\lambda$ combinations is due different $\Delta\Phi$ angles.
@@
In Figure 19 the parameter $p$ found is presented for all the simulations, in respect to $|\lambda|$ in $x$ axis (different markers for $\lambda>0$ and $\lambda<0$), the angle $\alpha_{1}$ (of the dominating field) is indicated by the point color and angle $\alpha_{2}$ by the point size. The figure demonstrates the significant variability in $p$ and shows a clear correlation between its average value and $|\lambda|$. The average value of $p$ is increased with $\lambda$ up to $|\lambda|=100.$ Note here that the high value of $p$ does not mean that, in this case, the system produces more flux but that the flux increases faster.
@@
Additionally, as $|\lambda|$ increases, the effect of the less magnetized star on the magnetosphere structure should eventually be saturated. In this direction simulations for $|\lambda|=\sqrt{10^{3}},\sqrt{10^{5}},10^{3}$ are conducted and verify, that beyond $|\lambda|=100$ the value of $p$ stops increasing.
@@
On the other hand, the dependence of $p$ on the magnetic moment orientation is not that clear. Although for $|\lambda|\geq 10$, there seems to be a tendency for $p$ to be high for cases where the dominant star has an inclination angle $\alpha_{1}=45^{\circ}$ and less for cases with $\alpha_{1}=90^{\circ}$ this is not always consistent as the specific orientation of the magnetic moments is important and the power law lacks the ability to describe accurately the periodic variability in the flux. Note that in the Figure  19 for $|\lambda|=1,10$ we have multiple points with the same combinations of $\alpha_{1}$ and $\alpha_{2}$ due to the different $\Delta\Phi$ these cases correspond to.

### v2

@@ 3.1.1 Aligned and Antialigned Magnetic Moments
Figure 6 presents the distributions of the Poynting flux on a spherical surface with a radius of $13r_{\star}$ for the aligned $\uparrow\uparrow$ (left-hand column) and antialigned $\uparrow\downarrow$ (right-hand column) cases at the indicated snapshots. While the sky distributions are qualitatively similar, the magnitude of the total flux at a given time (in agreement with C. Palenzuela et al., 2013a) is higher in the antialigned case compared to the aligned case. In particular, we find it to be approximately 4 times higher than the aligned case. The time evolution of the total flux for both cases is shown in Figure 7.
This difference arises from the magnetospheric solution realized by the system, as discussed earlier. When the magnetic moments of the two stars are aligned $\uparrow\uparrow$, the magnetic field lines originating closer to the poles of the stars diverge from the orbital axis toward the equator before curving back to it with some of them connecting with the opposite magnetic pole of the same star. A detailed evolution of the magnetic field lines during the evolution, for a similar configuration, is described in J. F. Mahlmann & A. M. Beloborodov (2025). This structure diminishes the time-averaged open flux $\Phi_{\rm open}$, thereby reducing the emitted EM power. Additionally, it produces a flux distribution with the dark zones, similar to that shown in Figure 6 (and similar to Figure 13 presenting the Poynting flux distribution in C. Palenzuela et al., 2013a), corresponding to the interface separating the two interacting magnetospheres.
@@
On the other hand, antialigned magnetic moments $\uparrow\downarrow$ lead to a significantly different global solution with higher Poynting flux. This is somewhat surprising, especially considering that, in the vacuum case, antialigned moments result in partial cancellation of dipolar components in favor of higher multipoles (also see C. Palenzuela et al., 2013a). Instead, in this scenario, magnetic field lines from the inner hemispheres of each star (the sides facing one another) connect directly between the stars. Some field lines originating from the outer hemispheres curve toward the magnetic poles of the other star and connect to the outer hemisphere of that star. However, field lines that originate closer to the equator tend to be longer, and those originating at sufficiently high polar angles may not curve inward enough to connect to the second star before crossing the equator. Once they cross the equator, they no longer connect with the other star, as they encounter the pole that has the same polarity. Consequently, these lines remain open and escape to infinity. This magnetosphere structure results in a significantly higher $\Phi_{\rm open}$ than the aligned case and a Poynting flux distribution without the dark zone observed in the aligned case. It is similar to the one reported in E. R. Most & A. A. Philippov (2022).
Figure 9: The polar caps, i.e., instantaneous zones of open flux, of the two stars in a Mollweide projection in the case with $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ},\,\lambda=-1$, and $\Delta\Phi=0^{\circ}$(
@@ 3.1.2 Inclined Magnetic Moments
Inclined magnetic moments introduce significant richness into the solutions, breaking both the latitudinal symmetry and the reflection symmetry of the emergent Poynting flux across the axial plane that passes through the center of the system, perpendicular to the line connecting the two stars. Figure 8 depicts configurations where the stars have non-zero-inclination angles for both the quasi-aligned ($\lambda>0$, first and second columns) and quasi-antialigned ($\lambda<0$, third and fourth columns) solutions. In the first ( $\uparrow$ $\uparrow$ ) and third columns ( $\uparrow$ $\downarrow$ ), a star has an inclination angle of $45^{\circ}$, whereas, in the second ( $\uparrow$ $\uparrow$ ) and fourth columns ( $\uparrow$ $\downarrow$ ), it has an inclination angle of $90^{\circ}$. Notably, the second and fourth columns correspond to configurations that are mirror images of each other under a rotation by $\pi$ around the axis, defined by the line connecting the two stars. Additionally, in the corotating frame, the two stars appear to rotate because they are irrotational in the observer frame.
Unlike the fully aligned or antialigned cases (§3.1.1), the presence of inclined magnetic moments introduces a more pronounced impact of orbital motion on the magnetosphere’s structure. The resulting magnetospheres evolve dynamically, exhibiting periodic variations and eruptions of Poynting flux linked to the orbital phases. As the stars orbit each other, magnetic fields originating from one star are carried along with its motion, anchored on its surface. As a result, the system reorganizes during a single rotation (see the middle and bottom rows in Figure  5). Field lines can dynamically evolve throughout the inspiral; some may connect one star to the other, while others that initially link the two stars may break and reconfigure. As a result, field lines anchored to a star’s surface can transition between being open, closed, or connecting with the companion star.
Thus, the polar caps, defined as footpoints of field lines extending to infinity (i.e., open flux), evolve over time, as shown in Figure 9. This evolution follows the orbital periodicity, with the variability becoming increasingly pronounced as the separation between the stars decreases during inspiral. The polar cap regions are found by integrating magnetic field lines using a fourth-order Runge-Kutta method, classifying as open those that cross a cylinder of radius $25r_{\star}$ (if sufficiently large, the value of this cylindrical radius does not seem to impact significantly the size and shape of the polar caps; see also N. Ortiz et al., 2022) and height of $50r_{\star}$. In Figure 9, the polar caps of both stars are shown at two snapshots separated by half a rotation, for the case with $\lambda=-1$, $\alpha_{1}=0^{\circ}$, and $\alpha_{2}=45^{\circ}$ ( $\uparrow$ $\downarrow$ ).
@@
Figure 10 shows the flux distributions on a spherical surface with radius $r=13r_{\star}$ corresponding to the magnetospheric structures and snapshots shown in Figure 8. Consistent with the discussion above, unlike the zero-inclination cases discussed previously, the flux distribution in these cases is no longer symmetric between the northern and southern hemispheres. While the flux pattern may appear symmetric over a full orbital period, the instantaneous flux evolves throughout the inspiral, leading to unequal fluxes in the two hemispheres at any given moment (E. R. Most & A. A. Philippov, 2022). This asymmetry in the instantaneous and time-averaged Poynting flux pattern indicates the presence of EM forces and torques acting on the system, influencing the motion of the NSs and potentially affecting the postmerger remnant. This aspect is further explored in Section 3.2.
@@
The variability and dependence on magnetic configuration of the EM flux are also evident in Figure 12, which shows the total flux passing through the spherical surface with $r=13r_{\star}$ as a function of time. For $\lambda>0$ (left panel), the total flux increases when the stars have non-zero-inclination angles and rises further as the angles increase. In contrast, for $\lambda<0$ (right panel), the total flux decreases with increasing inclination angles. Notably, when both stars have inclination angles of $90^{\circ}$ and $\lambda<0$, the total flux can drop to nearly zero at specific times during the evolution. This is illustrated in the bottom row of Figure  5, where two snapshots are shown, separated by an orbital phase of $\pi/2$. When the magnetic poles of the two stars with the same polarity face each other (left panel), field lines bend away, twist, and predominantly connect each star’s opposite magnetic poles, allowing a modest number of field lines to remain open. On the other hand, after a time evolution corresponding to the rotation of $\pi/2$ (right panel), the magnetic moments are oriented side by side, pointing in opposite directions. This configuration leads to even fewer open field lines and a significantly lower escaping flux compared to the first snapshot.
This behavior reflects the fact that the aligned and antialigned configurations represent two extreme cases of the angle between the magnetic moments, directly influencing the degree of connectivity between the stars and the amount of open flux $\Phi_{\rm open}$.
For completeness, it is worth noting how the above results change in the opposite limit of tidally locked (synchronously rotating) stars, where $\omega_{1}=\omega_{2}=\Omega$. The differences relative to the irrotational cases depend on the magnetic configuration. In the synchronized and aligned case, the overall magnetospheric structure and resulting Poynting flux would be nearly identical to the aligned irrotational configuration. This is because the relative orientation between the magnetic moments and the orbital motion remains constant, resulting in minimal time variation in the magnetosphere. In the synchronized and antialigned configuration, some magnetic field lines connect the two stars, reducing the number of open field lines available to transport the EM energy. This leads to a lower Poynting flux compared to the corresponding aligned case, in contrast to what was found in the irrotational scenario. However, for cases where at least one of the stars has an inclined magnetic moment, the differences are more pronounced. The synchronization prevents the time-varying reconfiguration of field lines seen in the irrotational cases. As a result, the Poynting flux evolution arises solely from the shrinking orbital separation, which increases the polar cap area and flux magnitude.
@@ 3.1.3 Unequal Magnetization Configurations
Figure 12: The total Poynting flux crossing a spherical surface with radius $13r_{\star}$ for various inclination angle combinations of the two stars, shown for cases with $\lambda=1$ (left) and $\lambda=-1$ (right). The dotted line, in both plots, is the orbital average Poynting flux $\overline{\langle L\rangle}$ derived from the scaling law of Eq. 15 for $|\lambda|=1$.
@@
Figure 13: The magnetospheric structure at $\tau=-4.27$ ms, similar to Figure 8, is shown with increasing absolute values of the magnetic moment ratio from top to bottom. The left column illustrates the impact of the magnetic moment ratio in the antialigned case ($\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$) (
@@
The distribution of orbit-averaged emission in cases with $|\lambda|>1$ becomes increasingly concentrated around the equator (see also Figures 13 and 4 in C. Palenzuela et al., 2013a; E. R. Most & A. A. Philippov, 2022, respectively), as shown in the left column of Figure 14. As $|\lambda|$ increases, the emission progressively resembles that of an isolated pulsar with the same inclination angle as the dominant star, similar to Fig. 2.
@@ 3.1.4 Anisotropic Emission Patterns
Figure 14: The distribution of the Poynting flux across a spherical surface of radius $13r_{\star}$, integrated over the last 7.7 ms of the inspiral. White lines indicate the areas containing $50\%$ and $90\%$ of the total (integrated over the entire surface) flux. (Left) The two stars have antialigned magnetic moments $\uparrow\downarrow$, with magnetic moment ratios $\lambda=-1,-10,-100$ from top to bottom, respectively. (Right) Cases with $\lambda=-10$ and $\alpha_{1}=45^{\circ},\alpha_{2}=0^{\circ}$ (
@@
In particular, in general, $50\%$ of the escaping EM energy is in $50\%$ of the total sky area, while, in some configurations, it is directed toward $15\%$ of the sky (see Figure 16).
In specific configurations, the emission anisotropy is more uniform, similar to an isolated pulsar. In these cases, the time-averaged energy distribution is independent of the azimuthal angle; thus, potential observers located at different azimuths would receive similar amounts of fluence with the observer’s polar angle, $\zeta$, being the only determining parameter.
Figure 15: The total Poynting flux crossing a spherical surface radius $13r_{\star}$ throughout the system’s evolution for cases with $\alpha_{1}=\alpha_{2}=0$ and $\lambda<0$ ($\uparrow\downarrow$) and progressively increasing $|\lambda|$ values.
@@ 3.1.5 Scaling Laws for Poynting Luminosity and the Binary Braking Index
Simulating the evolution of the magnetosphere for a variety of configurations allows us to investigate how the escaping EM flux depends on the orbital frequency, $\Omega$, and consequently on time via Eq. 5. The simplest way to quantify this dependency is to assume a power-law relation between the luminosity and the orbital frequency analogous to an isolated pulsar, such that $\langle L\rangle\propto\Omega^{p}$, where $p$ is the power-law index. The classical dipolar isolated pulsar magnetosphere in either vacuum or FF regimes follows $p\equiv 4$ for the far-field, and it has often been assumed by heuristic arguments arguments that this law roughly (e.g., $p\sim 3.3-4.7$; see discussion) holds in the double NS case. By fitting this function to the evolving total Poynting flux from each BNS simulation, we obtain the distribution of the values of the exponent $p$, shown in Figure 17.
@@
Figure 17: Histogram of the power-law index $p$ obtained from fitting of the relation $\langle L\rangle\propto\Omega^{p}$ shown in gray for all data points and in red only for the peak values. The distribution of $p$ spans a wide range from $\sim 1$ to >6>6, in contrast to the $p\equiv 4$ value for an isolated dipolar pulsar.
@@
Notably, the parameter $p$ now spans a broad range of values spanning from 1 to more than 6. While there are many cases around $p=4$, the binary system shows cases where $p$ deviates significantly, either higher or lower, due to the complex interactions between the two stars’ magnetospheres. This surprisingly large deviation from the conventional dipole braking index $n\equiv p-1=3$ highlights the richness of near-field effects permitted in the double NS case that modify far-fields and escape radiation.
@@
Figure 18: Examples fit of the Poynting flux luminosity with a power law $\langle L\rangle\propto\Omega^{p}$. Top row: (left) the case with $\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$, and $\lambda=-1$. (Right) The case with $\alpha_{1}=45^{\circ},\alpha_{2}=0^{\circ}$, and $\lambda=10$ demonstrates that, although the power law cannot describe the periodic fluctuations, it captures well the trend of the luminosity. Bottom row: (left) the case with $\alpha_{1}=90^{\circ},\alpha_{2}=90^{\circ},\lambda=-1$, and $\Delta\Phi=0^{\circ}$ highlights that, although the flux can vary significantly, dropping to very low values and affecting the flux dependence on $\Omega$ the power law still describes quite well the averaged trend. (Right) The case with $\alpha_{1}=0^{\circ},\alpha_{2}=0^{\circ}$, and $\lambda=100$, which consists of an example of not that accurate description of the Poynting flux luminosity’s evolution with the power law. In all the cases, the dotted line indicates the moving average for comparison.
@@
Figure  18 showcases characteristic examples of the Poynting flux luminosity fits. While the power-law model captures the gross secular orbit-averaged trend of the luminosity, it does not account for the periodic fluctuations around the fitted $L(\Omega)$ values. This is particularly evident for high inclination angles (bottom left panel in Figure 18), where, as described in the previous section, the total flux fluctuates dramatically, occasionally dropping to near-zero levels. This extreme variability weakens the overall dependence of the flux on $\Omega$, causing the value of $p$ to decrease, often approaching values close to 1, as seen in Figure 17. Importantly, a fitted $p\sim 1$ does not necessarily imply an induced split-monopolar or wind configuration (e.g., A. K. Harding et al., 1999; N. Bucciantini et al., 2006). Instead, it reflects the presence of intense, episodic bursts of Poynting flux that produce similar luminosities from one orbit to the next.
@@
In some cases, the evolution of the Poynting flux appears to transition between two distinct regimes (bottom right panel in Figure 18), resulting in a fitted $p$ that represents an average rather than a single consistent trend. For instance, in the configuration shown in the bottom right panel, where both magnetic moments are aligned (to each other and to $\boldsymbol{\Omega}$), and one star’s field dominates, no field lines connect the two stars. At first, only the region of the open field lines of the primary star is influenced mostly by the secondary, with some bending around the companion and closing back onto the primary. This phase corresponds to the initial rise behavior in Poynting flux. However, as the inspiral progresses, the companion moves into the closed field region of the primary, leading to a transition where more field lines open. Since no lines can directly connect the two stars in this scenario, these newly opened field lines increase the Poynting flux faster, modifying its overall evolution.
@@
Despite these variations, we find that, in the majority of cases, the average Poynting flux evolution follows a single well-defined power-law trend over the fitting interval in our simulations. Instances where the flux transitions between two regimes, as seen in the bottom right panel of Figure 18, remain relatively rare.
@@
Figure 19: Scatter plot showing the power $p$ derived from the power-law fitting $\langle L\rangle\propto\Omega^{p}$ for the different simulated cases. On $x$-axis is $\lambda$ in logarithmic scale while the color and the size of each point indicate the two inclination angles $\alpha_{1}$ and $\alpha_{2}$ according to the legends. The circles are cases with $\lambda>0$, while the triangles are with $\lambda<0$. The different values of the exponent $p$ for the same $\alpha_{1}$, $\alpha_{2}$ and $\lambda$ combinations is due different $\Delta\Phi$ angles.
@@
In Figure 19 the parameter $p$ found is presented for all the simulations; in respect to $|\lambda|$ in $x$-axis (different markers for $\lambda>0$ and $\lambda<0$), the angle $\alpha_{1}$ (of the dominating field) is indicated by the point color, and the angle $\alpha_{2}$ is indicated by the point size. The figure demonstrates the significant variability in $p$ and shows a clear correlation between its average value and $|\lambda|$. The average value of $p$ is increased with $\lambda$ up to $|\lambda|=100.$ Note here that the high value of $p$ does not mean that, in this case, the system produces more flux, but it means the flux increases faster.
@@
Additionally, as $|\lambda|$ increases, the effect of the less magnetized star on the magnetosphere structure should eventually be saturated. In this direction, simulations for $|\lambda|=\sqrt{10^{3}},\sqrt{10^{5}},10^{3}$ are conducted and verify that beyond $|\lambda|=100$ the value of $p$ stops increasing.
@@
On the other hand, the dependence of $p$ on the magnetic moment orientation is not that clear. Although for $|\lambda|\geq 10$, there seems to be a tendency for $p$ to be high for cases where the dominant star has an inclination angle $\alpha_{1}=45^{\circ}$ and less for cases with $\alpha_{1}=90^{\circ}$; this is not always consistent as the specific orientation of the magnetic moments is important, and the power law lacks the ability to describe accurately the periodic variability in the flux. Note that in Figure  19 for $|\lambda|=1,10$ we have multiple points with the same combinations of $\alpha_{1}$ and $\alpha_{2}$ due to the different $\Delta\Phi$ these cases correspond to.